Problem 2-18 (LG 2-8)

You note the following yield curve in The Wall Street Journal. According to the unbiased expectations theory, what is the one-year forward rate for the period beginning two years from today, 3f1? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

Maturity Yield
One day 2.45 %
One year 2.57
Two years 2.81
Three years 2.92
Answer #1

Forward rate for 1 year, commencing 2 years from now is calculated using the following formula:

3f1 = [(1+s3)^3/ (1+s2)^2]-1

Where s2= spot rate for 2 years and s3= spot rate for 3 years

Given, s2= 2.81% and s3= 2.92%

Therefore, one year forward rate for the period beginning 2 years from today= ((1+2.92%)^3/(1+2.81%)^2)-1

=(1.0292^3/1.0281^2) -1.

=(1.090183/1.05699)-1 = 1.031404-1 = 3.14%